### v1

@@ Acknowledgements.
Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic during this work are greatly appreciated. This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance Award No. DE-SC0013617 and under the STREAMLINE Collaboration Award No. DE-SC0024586.
Appendix A A heuristic proof for the real spectrum of a specific type of CEE
If the training calculations fulfill the requirements mentioned at the end of Sec. III.2, the emulator components have the same generic structure. For example,
$\begin{bmatrix}H\end{bmatrix}=\begin{pmatrix}M_{cs},M_{hm}\\ M_{hm}^{\ast},M_{cs}^{\ast}\end{pmatrix}$ (60)
with $M_{cs}$ as a $N_{b}\times N_{b}$ complex symmetric matrix and $M_{hm}$ a Hermitian matrix of the same size. The basis choice for this representation is: $\{|\Psi^{tr}_{1}\rangle,,,|\Psi^{tr}_{N_{b}}\rangle,|\tilde{\Psi}^{tr}_{1}\rangle,,,|\tilde{\Psi}^{tr}_{N_{b}}\rangle\}$ and $\{\langle\tilde{\Psi}^{tr}_{1}|,,,\langle\tilde{\Psi}^{tr}_{N_{b}}|,\langle\Psi^{tr}_{1}|,,,\langle\Psi^{tr}_{N_{b}}|\}$. Since $M_{hm}^{\ast}=M_{hm}^{T}$, $\begin{bmatrix}H\end{bmatrix}$ is a complex symmetric matrix. It can then be shown that
$\eta\begin{bmatrix}H\end{bmatrix}$ $\equiv\begin{bmatrix}\overline{H}\end{bmatrix}=\begin{pmatrix}M_{hm}^{\ast},M_{cs}^{\ast}\\ M_{cs},M_{hm}\end{pmatrix}\ ,\ \text{with}$ (61)
$\eta$ $\equiv\begin{pmatrix}0,1\\ 1,0\\ \end{pmatrix}\ .$ (62)
Now $\begin{bmatrix}\overline{H}\end{bmatrix}$ is Hermitian, meaning
$\begin{bmatrix}\overline{H}\end{bmatrix}^{\dagger}$ $=\begin{bmatrix}\overline{H}\end{bmatrix}\ ,$ (63)
$\eta^{-1}\begin{bmatrix}H\end{bmatrix}^{\dagger}\eta$ $=\begin{bmatrix}H\end{bmatrix}\ .$ (64)
The last equation shows that $\begin{bmatrix}H\end{bmatrix}$ is $\eta$-pseudo-Hermtian Mostafazadeh (2002a, b). By the same argument, the norm matrix also satisfy
$\eta^{-1}\begin{bmatrix}N\end{bmatrix}^{\dagger}\eta$ $=\begin{bmatrix}N\end{bmatrix}\ .$ (65)
The generalized eigenvalue problem can be transformed as follows.
$\begin{bmatrix}H\end{bmatrix}v$ $=$ $E\begin{bmatrix}N\end{bmatrix}v$ (66)
$\eta\begin{bmatrix}H\end{bmatrix}\eta^{-1}\eta v$ $=$ $E\eta\begin{bmatrix}N\end{bmatrix}\eta^{-1}\eta v$ (67)
$\begin{bmatrix}H\end{bmatrix}^{\dagger}\eta v$ $=$ $E\begin{bmatrix}N\end{bmatrix}^{\dagger}\eta v$ (68)
$\begin{bmatrix}H\end{bmatrix}^{\ast}\eta v$ $=$ $E\begin{bmatrix}N\end{bmatrix}^{\ast}\eta v$ (69)
$\begin{bmatrix}H\end{bmatrix}(\eta v)^{\ast}$ $=$ $E^{\ast}\begin{bmatrix}N\end{bmatrix}(\eta v)^{\ast}$ (70)
The second before the last step invokes that $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ are symmetric. Therefore, $E$ and $E^{\ast}$ are both the eigenvalues, which generalizes the same conclusion of Ref. Mostafazadeh (2002a) about $\eta$-pseudo-Hermtian $\begin{bmatrix}H\end{bmatrix}$ but with identity norm matrix.
This reasoning narrows the eigenvalues (or locations of the BCT poles) to be on the real axis or distributed symmetrically on the two sides of the real axis. As argued in Sec. II.3, the approximated $\mathcal{A}$ with poles away from the real axis should have larger errors than the approximation with all poles on the real axis, considering the training calculations are distributed symmetrically on both sides of the real axis. The effective RBM subspace projection should produce an optimal approximation, eliminating the existence of the complex conjugate pairs in the eigenvalues.
@@ Appendix B Two-body
For the $s$-wave channel, we employ a Gaussian form factor,
@@
For the $p$-wave case, we have
@@
To get the $T$-matrix in the 2nd Riemann sheet (with the positive real-$E$ axis defined as the branch cut), we must use the correct branch for the $\sqrt{D}$ value.
@@ C.1 Analytical formulas and equations
We first introduce important variables, such as $Z_{12}$ and $Z_{14}$, which will appear in the following formulas and equations. Here, $Z$s are defined as [^23] Here, the orderings of the subscripts in $Z_{12}$ and $Z_{14}$ are not relevant. Their definitions differ from those in a previous work Zhang and Furnstahl (2022) by factors of $\xi^{2}$ and proper integer powers of $\sqrt{4\pi}$. This difference is due to the partial wave basis implemented here, while in Ref. Zhang and Furnstahl (2022), they were defined as averaged over the solid angle of the momentum vector.
$Z_{21}({P}_{2},{P}_{1},E)=\langle{P}_{2},g|G_{0}(E)|{P}_{1},g\rangle=(-)(2\pi)\frac{M}{P_{1}P_{2}}\left(\frac{C}{\Lambda_{2}}\right)^{2}e^{-\frac{5\left(P_{1}^{2}+P_{2}^{2}\right)}{8\Lambda_{2}^{2}}}\left\{e^{\tilde{D}}\left[\Gamma\left(0,\tilde{D}-\tilde{d}\right)-\Gamma\left(0,\tilde{D}+\tilde{d}\right)\right]\right\}\ ,$ (80)
@@
[…]
Here, $\xi$ is the dimer bound state normalization parameter in the $|\phi_{1}\rangle\equiv|P_{rel}\rangle|\varphi_{B}\rangle$ (see Eq. (75)).
@@
Note the sub and super-scripts $s$ are intended to explicitly show how the sources depend on the couplings and $E_{rel}^{s}\equiv P_{rel}^{2}/(2\mu^{1})-B_{2}^{s}$ depending on the dimer binding energy $B_{2}^{s}$. In this work, we will sometimes fix the parameters inside the source, whose values can differ from those in the $H$.
@@
[…]
@@
Here $\Theta(a\leq x\leq b)$ is $1$ when $x\in[a,b]$, and 00 otherwise.
In our three-body model, we can take advantage of the potentials’ separable nature to reduce the dimensionality of the Faddeev equations. Let’s define
@@
To solve $|\psi_{1}\rangle$, based on Eq. (53), we can express $|\psi_{4}\rangle$ in terms of $|\psi_{1}\rangle$:
@@
[…]
@@
Note $\Gamma\left(0,-tE^{+}\right)=E_{1}(-tE^{+})$ (DLMF, , Eq. 8.19.1) which cares about the $i0^{+}$ in $E^{+}\equiv E+i0^{+}$.
Considering that in Eq. (94), the three terms in the right side are proportional to $G_{0}t_{1}$, we introduce a new state $|F_{1}\rangle$,
@@
Eq. (96) suggests we only need to know $\langle P,g|F_{1}\rangle$ to compue $|\psi_{1}\rangle$. We first define $\ F_{1}(P)\equiv\langle P,{g}|F_{1}\rangle$, and then express $|\psi_{1}\rangle$ in terms of $F_{1}$:
@@
[…]
@@
The inhomogeneous equation (94), which has two variables $P$ and $q$, now reduces to Eq. (99) with just one variable $P$. Interestingly, the equation for $F_{1}(P)$ is similar to the equation for scattering amplitude in Ref. Zhang and Furnstahl (2022), but $E$ could be complex here.
@@
Plugging this expression back into Eq. (58), we get
[…]
@@
Eq. (105) shows that if we know $\langle P,{g}|\tilde{\psi}_{1}\rangle\equiv\tilde{F}_{1}(P)$, we can plug it on the right side and get $\langle P,q|\tilde{\psi}_{1}\rangle$. Therefore, let’s first look at the equation for $\tilde{F}_{1}(P)$:
[…]
Knowing $\tilde{F}_{1}(P)$, we can compute
@@
[…]
@@
In these integral equations, we need to have the explicit expression for the kernels, including $\langle P,{g}|G_{1}^{\ast}\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)|P^{\prime},g\rangle$ and $\langle P,q|G_{1}^{\ast}\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)|P^{\prime},g\rangle$. By expanding $G_{1}=G_{0}+G_{0}t_{1}G_{0}$, we get
@@
$=\frac{1}{E^{\ast}-\frac{P^{2}}{2\mu^{1}}-\frac{q^{2}}{2\mu_{1}}}\bigg{(}\langle Pq|\mathbb{P}|P^{\prime}{g}\rangle+3\tau_{4}^{\ast}g_{4}(P,q)Z_{14}(P^{\prime},E^{\ast})+{g}(q){\tau}_{1}(E^{\ast}-\frac{P^{2}}{2\mu^{1}})2\bar{Z}^{00}(P,P^{\prime},E^{\ast},E^{\ast},1,1)\bigg{)}$ (111)
In the derivations, we make use of the fact that
@@
We also need to know the driving terms explicitly in the integral equations. First,
@@
[…]
@@
[…]
@@
As a partial consistency check of the formulas derived so far, we note first that $\xi F_{1}(P_{rel})$, as the solution of Eq. (99) with $E=E_{0}$ and $E_{0}\equiv E_{rel}-B_{2}+i0^{+}$, is the full on-shell scattering amplitude as defined according to Eq. (56). To prove this, we go back to Eq. (56) and make the following derivations:
@@
$=\langle\phi_{1}(E_{0})|\big{[}(E_{0}-H_{1})G_{1}(E_{0})V_{1}|F_{1}(E_{0})\rangle\big{]}=\xi\langle P_{rel},g|G_{0}(E_{0})V_{1}|F_{1}(E_{0})\rangle$ (125)
$=\xi\lambda\langle g|G_{2b,0}(-B_{2})|g\rangle F_{1}(P_{rel})=\xi F_{1}(P_{rel})$ (126)
The first step comes from that $(H_{0}+V_{1})|\phi_{1}(E_{0})\rangle=E_{0}|\phi_{1}(E_{0})\rangle$. From Eq. (122) to (123), we utilize the definition of the Faddeev components in Eq. (47). The step from Eq. (123) to (124) is based on Eq. (50). Moreover, in Eq. (124), we would conclude that the matrix element becomes zero if we apply $H_{1}$ to the left. However, $H_{1}$ can not be applied to $\langle\phi_{1}|$ via integration by parts because both $\langle\phi_{1}|$ and $|\psi_{1}\rangle$ have oscillating asymptotics in the coordinate space when particle-1 is far away from dimer-23. Finally, in the last step, we use the fact that $\langle g|G_{2b,0}(-B_{2})|g\rangle=1/\lambda$ because of the bound-state pole in the $\tau$ function.
Meanwhile, the other route to verify $\xi F_{1}(P_{rel})$ as the on-shell scattering amplitude, when $E=E_{0}$, is to inspect its Eq. (99) directly. It is easy to see the equation is the same as the one for on-shell scattering amplitude computed in Ref. Zhang and Furnstahl (2022). Therefore, we see the consistency of the two routes and, thus, the consistency of the formula derivation in this section.

### v2

@@ Acknowledgements.
Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic during this work are greatly appreciated. The author also thanks Kevin Fossez, Sebastian König, Simin Wang and Witold Nazarewicz for valuable comments on the manuscript. This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance Award No. DE-SC0013617 and under the STREAMLINE Collaboration Award No. DE-SC0024586.
Appendix A On the Hermiticity of the generalized eigenvalue problem in a special type of CEE
If the training calculations fulfill the requirements mentioned at the end of Sec. III.2, i.e., the source states $|{S}_{\alpha}\rangle=|\tilde{{S}}_{\alpha}\rangle$ and are invariant under time-reversal transformation, and both $E_{\alpha}^{\mathrm{tr}}$ and its complex conjugate $E_{\alpha}^{\mathrm{tr}\ast}$ are included in the training set, $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ in fact have the same generic structure and are Hermitian. For example,
$\begin{bmatrix}N\end{bmatrix}=\begin{pmatrix}M_{hm},M_{cs}\\ M_{cs}^{\ast},M_{hm}^{\ast}\end{pmatrix}\$, (64)
with $M_{cs}$ as a $N_{b}\times N_{b}$ complex symmetric matrix and $M_{hm}$ a Hermitian matrix of the same size.
The choice of the basis for this representation is
$\bigg\{$ $|\Psi(E^{\mathrm{tr}}_{1},\bm{\theta}^{\mathrm{tr}}_{1})\rangle,,,|\Psi(E^{\mathrm{tr}}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})\rangle,$
$|\Psi(E^{\mathrm{tr}\ast}_{1},,\bm{\theta}^{\mathrm{tr}}_{1})\rangle,,,|\Psi(E^{\mathrm{tr}\ast}_{N_{b}},,\bm{\theta}^{\mathrm{tr}}_{N_{b}})\rangle\bigg\}$ (65)
and
$\bigg\{$ $\langle\tilde{\Psi}(E^{\mathrm{tr}\ast}_{1},\bm{\theta}^{\mathrm{tr}}_{1})|,,,\langle\tilde{\Psi}(E^{\mathrm{tr}\ast}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})|,$
$\langle\tilde{\Psi}(E^{\mathrm{tr}}_{1},\bm{\theta}^{\mathrm{tr}}_{1})|,,,\langle\tilde{\Psi}(E^{\mathrm{tr}}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})|\bigg\}$ (66)
To see the Hermiticity, one should note that with $|{S}(\bm{\theta})\rangle=|\tilde{{S}}(\bm{\theta})\rangle$, the solutions of Eqs. (31) and (32) have the following connections:
$|\tilde{\Psi}(E,\bm{\theta})\rangle$ $=|\Psi(E^{\ast},\bm{\theta})\rangle\ ,$ (67)
$|\Psi(E,\bm{\theta})\rangle$ $=|\tilde{\Psi}(E^{\ast},\bm{\theta})\rangle\ ,$ (68)
as they are analytical functions of $E$ away from the real-$E$ axis. Therefore, the basis sets are, in fact, the same as
$\bigg\{$ $|\Psi(E^{\mathrm{tr}}_{1},\bm{\theta}^{\mathrm{tr}}_{1})\rangle,,,|\Psi(E^{\mathrm{tr}}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})\rangle,$
$|\tilde{\Psi}(E^{\mathrm{tr}}_{1},,\bm{\theta}^{\mathrm{tr}}_{1})\rangle,,,|\tilde{\Psi}(E^{\mathrm{tr}}_{N_{b}},,\bm{\theta}^{\mathrm{tr}}_{N_{b}})\rangle\bigg\}$ (69)
and
$\bigg\{$ $\langle\Psi(E^{\mathrm{tr}}_{1},\bm{\theta}^{\mathrm{tr}}_{1})|,,,\langle\Psi(E^{\mathrm{tr}}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})|,$
$\langle\tilde{\Psi}(E^{\mathrm{tr}}_{1},\bm{\theta}^{\mathrm{tr}}_{1})|,,,\langle\tilde{\Psi}(E^{\mathrm{tr}}_{N_{b}},\bm{\theta}^{\mathrm{tr}}_{N_{b}})|\bigg\}\ .$ (70)
Indeed, both $\begin{bmatrix}H\end{bmatrix}$ and $\begin{bmatrix}N\end{bmatrix}$ are Hermitian.
One can also conclude that the diagonal blocks in Eq. (64) which are Hermitian are the complex conjugate of each other, considering that via Eqs. (33) and (34),
$\langle\tilde{\Psi}(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|\tilde{\Psi}(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})\rangle=\langle\mathbb{T}\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|\mathbb{T}\Psi(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})\rangle$
$=\langle\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|\Psi(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})\rangle^{\ast}\ .$ (71)
Similarly by using $\mathbb{T}$ transformation, I have
$\langle\tilde{\Psi}(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|\Psi(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})\rangle=\langle\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|\tilde{\Psi}(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})\rangle^{\ast}$
$=\langle\tilde{\Psi}(E_{\beta}^{\mathrm{tr}},\bm{\theta}_{\beta}^{\mathrm{tr}})|\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})\rangle\ ,$ (72)
i.e. the off-diagonal blocks in Eq. (64) are the complex conjugate of each other (from the first equal sign) and each block is complex symmetric (from the second equal sign). Based on the same argument, one can see that the same block structure for $\begin{bmatrix}H\end{bmatrix}$.
@@ Appendix B Two-body
For the $s$-wave channel, I employ a Gaussian form factor,
@@
For the $p$-wave case, I have
@@
To get the $T$-matrix in the 2nd Riemann sheet (with the positive real-$E$ axis defined as the branch cut), one must use the correct branch for the $\sqrt{D}$ value.
@@ C.1 Analytical formulas and equations
I first introduce important variables, such as $Z_{12}$ and $Z_{14}$, which will appear in the following formulas and equations. Here, $Z$s are defined as [^20] Here, the orderings of the subscripts in $Z_{12}$ and $Z_{14}$ are not relevant. Their definitions differ from those in a previous work [40] by factors of $\xi^{2}$ and proper integer powers of $\sqrt{4\pi}$. This difference is due to the partial wave basis implemented here, while in Ref. [40], they were defined as averaged over the solid angle of the momentum vector.
[…]
@@
$T_{\mathrm{Born}}=2\xi^{2}\lambda Z_{21}(P_{in},P_{in},E_{in})+3\lambda_{4}\xi^{2}Z_{41}(P_{in},E_{in})Z_{14}(P_{in},E_{in})+\int dPP^{2}\lambda\xi^{2}2Z_{12}(P_{in},P,E_{in})2Z_{21}(P_{in},P,E_{in})\ .$ (86)
Here, $\xi$ is the dimer bound state normalization parameter in the $|\phi_{1}\rangle\equiv|P_{\mathrm{rel}}\rangle|\varphi_{B}\rangle$ (see Eq. (77)).
@@
Note the sub and super-scripts $s$ are intended to explicitly show how the sources depend on the couplings and $E_{\mathrm{rel}}^{s}\equiv P_{\mathrm{rel}}^{2}/(2\mu^{1})-B_{2}^{s}$ depending on the dimer binding energy $B_{2}^{s}$. In this work, I will sometimes fix the parameters inside the source, whose values can differ from those in the $H$.
@@
$\langle P,q|\tilde{{S}}_{1}\rangle=\langle P,q|(1+\mathbb{P})V_{1}^{s}\mathbb{P}+3V_{4}^{s}|\phi_{1}\rangle=\langle P,q|{S}_{1}\rangle+\langle P,q|{S}_{4}\rangle+\lambda^{s}\xi_{s}\int dP^{\prime}P^{\prime 2}\langle P,q|\mathbb{P}|P^{\prime},{g}\rangle 2Z_{12}(P^{\prime},P_{in},E_{\mathrm{rel}}^{s})\ ,$ (89)
@@
Here $\Theta(a\leq x\leq b)$ is $1$ when $x\in[a,b]$, and 0 otherwise.
In the current three-body model, I take advantage of the potentials’ separable nature to reduce the dimensionality of the Faddeev equations. I first define
@@
To solve $|\psi_{1}\rangle$, based on Eq. (56), I can express $|\psi_{4}\rangle$ in terms of $|\psi_{1}\rangle$:
@@
Note, in the three-body Fock space, $V_{1}=\int dPP^{2}\lambda|P,{g}\rangle\langle P,{g}|$. Similarly, $t_{1}(E)=\int dPP^{2}\tau(e_{2})|P,{g}\rangle\langle P,{g}|$, with the full three-body total energy $E$ as its argument, $e_{2}=E-P^{2}/(2\mu^{1})$, and $P$ as the spectator’s momentum. $\tau(e_{2})$ is discussed in Appendix B. $t_{4}$ and the related $\tau_{4}$ have been discussed in Sec. V.1.1. Its analytical form is
@@
Note $\Gamma\left(0,-tE^{+}\right)=E_{1}(-tE^{+})$ [117, Eq. 8.19.1] which cares about the $i0^{+}$ in $E^{+}\equiv E+i0^{+}$.
Considering that in Eq. (96), the three terms in the right side are proportional to $G_{0}t_{1}$, I introduce a new state $|F_{1}\rangle$,
@@
Eq. (98) suggests that one only needs to know $\langle P,g|F_{1}\rangle$ to compute $|\psi_{1}\rangle$. I first define $F_{1}(P)\equiv\langle P,{g}|F_{1}\rangle$, and then express $|\psi_{1}\rangle$ in terms of $F_{1}$:
@@
$=\xi_{s}2\bar{Z}(P,P_{\mathrm{rel}},E_{\mathrm{rel}},E_{\mathrm{rel}},\frac{\lambda^{s}}{\lambda},\frac{\lambda_{4}^{s}}{\lambda_{4}})+\int dP^{\prime}P^{\prime 2}2\bar{Z}(P,P^{\prime},E,E,1,1){\tau_{1}}(E-P^{\prime 2}/(2\mu^{1}))F_{1}(P^{\prime})\ ,$ (101)
@@
The inhomogeneous equation (96), which has two variables $P$ and $q$, now reduces to Eq. (101) with just one variable $P$. Interestingly, the equation for $F_{1}(P)$ is similar to the equation for scattering amplitude in Ref. [40], but $E$ could be complex here.
@@
Plugging this expression back into Eq. (61), I get
$\left(E^{\ast}-H_{0}-V_{1}-\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)V_{1}\right)|\tilde{\psi}_{1}\rangle=|\tilde{{S}}_{1}\rangle+3V_{4}G_{4}^{\ast}|\tilde{{S}}_{4}\rangle\ \text{i.e.,}$ (106)
@@
Eq. (107) shows that if $\langle P,{g}|\tilde{\psi}_{1}\rangle\equiv\tilde{F}_{1}(P)$ is known, I can plug it on the right side and get $\langle P,q|\tilde{\psi}_{1}\rangle$. Therefore, one can first look at the equation for $\tilde{F}_{1}(P)$:
$\tilde{F}_{1}(P)=\langle P,g|G_{1}^{\ast}\left(|\tilde{{S}}_{1}\rangle+3t_{4}^{\ast}G_{0}^{\ast}|\tilde{{S}}_{4}\rangle\right)+\int dP^{\prime}P^{\prime 2}\langle P,{g}|G_{1}^{\ast}\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)|P^{\prime},g\rangle\lambda\tilde{F}_{1}(P^{\prime})\ .$ (108)
Knowing $\tilde{F}_{1}(P)$, I can compute
@@
$\langle P,q|\tilde{\psi}_{4}\rangle=\langle P,q|G_{4}^{\ast}\left[|\tilde{{S}}_{4}\rangle+V_{1}|\tilde{\psi}_{1}\rangle\right]\\ =\frac{1}{E^{\ast}-\frac{P^{2}}{2\mu^{1}}-\frac{q^{2}}{2\mu_{1}}}\bigg\{\langle P,q|\tilde{{S}}_{4}\rangle+g_{4}(P,q)\tau_{4}^{\ast}\langle g_{4}|G_{0}^{\ast}|\tilde{{S}}_{4}\rangle+\lambda g(q)\tilde{F}_{1}(P)+\lambda g_{4}(P,q)\tau_{4}^{\ast}\int dPP^{2}Z_{14}(P,E^{\ast})\tilde{F}_{1}(P)\bigg\}\ .$ (110)
@@
In these integral equations, I need to have the explicit expression for the kernels, including $\langle P,{g}|G_{1}^{\ast}\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)|P^{\prime},g\rangle$ and $\langle P,q|G_{1}^{\ast}\left(\mathbb{P}+3t_{4}^{\ast}G_{0}^{\ast}\right)|P^{\prime},g\rangle$. By expanding $G_{1}=G_{0}+G_{0}t_{1}G_{0}$, I get
@@
$=\frac{1}{E^{\ast}-\frac{P^{2}}{2\mu^{1}}-\frac{q^{2}}{2\mu_{1}}}\bigg(\langle Pq|\mathbb{P}|P^{\prime}{g}\rangle+3\tau_{4}^{\ast}g_{4}(P,q)Z_{14}(P^{\prime},E^{\ast})+{g}(q){\tau}_{1}(E^{\ast}-\frac{P^{2}}{2\mu^{1}})2\bar{Z}(P,P^{\prime},E^{\ast},E^{\ast},1,1)\bigg)$ (113)
In the derivations, I make use of the fact that
@@
I also need to know the driving terms explicitly in the integral equations. First,
@@
$=\frac{\tau_{1}(E^{\ast}-\frac{P^{2}}{2\mu^{1}})}{\lambda}\int dqq^{2}\frac{g(q)}{E^{\ast}-P^{2}/(2\mu^{1})-q^{2}/(2\mu_{1})}\langle P,q|\tilde{{S}}_{1}\rangle\ ,$ (116)
@@
$=3\tau_{4}^{\ast}\frac{\tau_{1}(E^{\ast}-\frac{P^{2}}{2\mu^{1}})}{\lambda}\left[\int dqq^{2}\frac{g(q)g_{4}(P,q)}{E^{\ast}-P^{2}/(2\mu^{1})-q^{2}/(2\mu_{1})}\right]\left[\int dPdqP^{2}q^{2}\frac{\langle P,q|\tilde{{S}}_{4}\rangle g_{4}(P,q)}{E^{\ast}-P^{2}/(2\mu^{1})-q^{2}/(2\mu_{1})}\right]$ (118)
@@
As a partial consistency check of the formulas derived so far, I note first that $\xi F_{1}(P_{\mathrm{rel}})$, as the solution of Eq. (101) with $E=E_{0}$ and $E_{0}\equiv E_{\mathrm{rel}}-B_{2}+i0^{+}$, is the full on-shell scattering amplitude as defined according to Eq. (59). To prove this, one goes back to Eq. (59) and makes the following derivations:
@@
$=\langle\phi_{1}(E_{0})|\big[(E_{0}-H_{1})G_{1}(E_{0})V_{1}|F_{1}(E_{0})\rangle\big]=\xi\langle P_{\mathrm{rel}},g|G_{0}(E_{0})V_{1}|F_{1}(E_{0})\rangle$ (127)
$=\xi\lambda\langle g|G_{2b,0}(-B_{2})|g\rangle F_{1}(P_{\mathrm{rel}})=\xi F_{1}(P_{\mathrm{rel}})$ (128)
The first step comes from that $(H_{0}+V_{1})|\phi_{1}(E_{0})\rangle=E_{0}|\phi_{1}(E_{0})\rangle$. From Eq. (124) to (125), I utilize the definition of the Faddeev components in Eq. (50). The step from Eq. (125) to (126) is based on Eq. (53). Moreover, in Eq. (126), one would conclude that the matrix element becomes zero if $H_{1}$ is applied to the left. However, $H_{1}$ can not be applied to $\langle\phi_{1}|$ via integration by parts because both $\langle\phi_{1}|$ and $|\psi_{1}\rangle$ have oscillating asymptotics in the coordinate space when particle-1 is far away from dimer-23. Finally, in the last step, I use the fact that $\langle g|G_{2b,0}(-B_{2})|g\rangle=1/\lambda$ because of the bound-state pole in the $\tau$ function.
Meanwhile, the other route to verify $\xi F_{1}(P_{\mathrm{rel}})$ as the on-shell scattering amplitude, when $E=E_{0}$, is to inspect its Eq. (101) directly. It is easy to see the equation is the same as the one for on-shell scattering amplitude computed in Ref. [40]. Therefore, one sees the consistency of the two routes and, thus, the consistency of the formula derivation in this section.